# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

### v1

@@ IV.2 Notes on the acceleration-convergence schemes
Now let us turn to the comparative analysis of the SC and two-potential schemes used to accelerate the convergence of the many-potential term in two gauges. In Fig. 8 (a) the ratio of the (quasi-) three-plus-potential to the many-potential contributions is plotted. In the small-$Z$ region the two-potential scheme in Coulomb gauge yields the smallest three-plus-potential contribution to be extrapolated. However, the accuracy there is limited by the remainder – the two-potential contribution itself. The non-monotonicity of the Coulomb gauge SC graph is explained by the fact that the quasi-three-plus potential contribution passes through zero and changes sign.
Figure 7: Comparison of two schemes in two gauges, extrapolated sum of the series (a) and leading power ($\sim\frac{1}{k^{2}}$) coefficient $C_{2}^{3+}$ (b) for (quasi-) three-plus-potential terms. Note the normalization to the many-potential contribution.
Let us discuss the extrapolation coefficients $C_{i}$ from Eq. (43). From our study it follows that their relative error grows quite rapidly with increasing $i$. Therefore we consider the first expansion coefficient $C_{2}$, plotted in Fig. 7 (b). This graph partially explains the convergence acceleration mechanism. In the area of small $Z$, the coefficient $C_{2}$ decreases by one or two orders of magnitude; in the region of large $Z$, it also decreases, but less effectively. Within the same gauge regardless of the tricks $C_{2}$ decreases almost identically illustrating the viability of the SC approximation.
Figure 8: Comparison of two schemes in two gauges, normalized PW expansion contributions. Data for $Z=18$ argon (a) and for $Z=92$ uranium (b) is shown.
We cannot conclude that the extraction of two-potential contributions reduces the minimum power in the expansion (43) to 3; we argue that the opposite situation occurs. The last statement has been verified empirically – the extrapolation procedure gave incorrect results. It can also be reasoned by the contributions graphs scaled by $k^{3}$, where the (quasi-)three-plus-potential contribution come out to a constant in all cases, see Fig. 8 (b). This graph also demonstrates that any acceleration scheme in the Coulomb gauge performs better than in the Feynman gauge, as the convergence to the asymptotics is much faster.
From the above analysis, it is difficult to conclude which specific acceleration scheme is better. However the SC scheme has a significant advantage: the ability to obtain a two-potential contribution without a PW expansion. In practice almost always the combination of the Coulomb gauge and the SC scheme allows one to obtain the most accurate self-energy correction for all $Z$.
@@ IV.3 Data: summary
The calculated SE corrections with the parameters from the Ref. hen are summarized in Tab. 2. In this table the zero-, one-, and most accurate many-potential contributions for three nuclei are listed in two gauges. One can see that the total uncertainty of the SE correction is mainly determined by the uncertainty of the many-potential contribution. Even taking into account the acceleration schemes. We present separately the many-potential, two-potential, and quasi-two-potential contributions in Tab. 3. Using this table one can compare the accuracy of the many-potential contribution obtained. Overall, it is clear that the acceleration tricks do reduce the extrapolation error and allow to obtain the many-potential contribution more accurately. In addition, there is a small but noticeable difference between our data and the data from Ref. hen.
Results with the parameters from the Ref. se_yerokh are presented in the similar manner in Tab. 5. An attractive feature of the Coulomb gauge can be seen from that table – the absence of the large reductions between the momentum (zero-, one-potential) contributions in the low-$Z$ region. It is clearly visible from the Fig. 9 (a) (compare the green and yellow curves). Various coordinate contributions are given in a separate table 6. In some cases we are able to reduce the numerical error compared with the work se_yerokh and the results obtained are in reasonable agreement with the corresponding work. Calculations for excited states using the Green’s function method are also performed, see Tab. LABEL:tab:we_yerokh_excited_states.
@@ V Conclusion
The self-energy correction for various states of hydrogen-like ions is calculated. Two gauges are considered, acceleration tricks are discussed and convenient formulas for the corresponding calculations are collected.
Overall, the total numerical uncertainty of the SE correction is mainly determined by the numerical uncertainty of the many-potential contribution. The contribution itself is significantly smaller in the Coulomb gauge than in the Feynman one. It could be seen, firstly, by the blue curves in the Fig. 9 (a), and secondly, in Fig. 9 (b), where the sum of the zero- and one-potential contributions is shown separately. Despite the smallness of the many-potential contribution, it does not allow obtaining a full self-energy correction with greater accuracy if no acceleration schemes are involved. An important feature of the Coulomb gauge SE calculations is the absence of the large reductions between the momentum (zero-, one-potential) contributions in the low-$Z$ region.
The acceleration schemes do improve the accuracy of the calculations performed. Using either the two-potential scheme or the SC scheme reduces the extrapolation error and allows one to obtain up to two new significant figures in the many-potential contribution. The combination of Coulomb gauge and SC scheme is most suitable (among those considered) for calculating the self-energy correction in hydrogen-like ions, especially when considering light nuclei. The next goal is to consider the Fried-Yennie gauge; we believe that the analysis performed will simplify the calculation of a wide range of more complex radiative corrections with the self-energy loop.

### v2

@@ IV.2 Application of the convergence-acceleration schemes
Now let us turn to the comparative analysis of the SC and two-potential schemes used to accelerate the convergence of the PW expansions for the many-potential contributions in the Feynman and Coulomb gauges. In Fig. 7(a), ratios of the three-plus-potential $\Delta\varepsilon_{a,x}^{\text{(3+)p}}$ and quasi-three-plus-potential $\Delta\tilde{\varepsilon}_{a,x}^{\text{(3+)p}}$ contributions to the total many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}$ are plotted. In the low-$Z$ region, the two-potential scheme in Coulomb gauge yields the smallest remainder which is to be extrapolated. However, in this case the overall accuracy is limited by the two-potential contribution, which also has to be treated within the PW expansion.
Figure 7: Comparison of two convergence-acceleration schemes in two gauges, Feynman and Coulomb ones: (a) three-plus-potential $\Delta\varepsilon_{a,x}^{\text{(3+)p}}$ and quasi-three-plus-potential $\Delta\tilde{\varepsilon}_{a,x}^{\text{(3+)p}}$ contributions normalized to the total many-potential contributions $\Delta\varepsilon_{a}^{\text{Mp}}$; (b) the same for the coefficients $C_{2}$ defined in Eq. (47).
Let us turn to the coefficients $C_{2}$ from Eq. (47) obtained for $\Delta{\varepsilon}_{a,x}^{\text{(3+)p}}$ and $\Delta\tilde{\varepsilon}_{a,x}^{\text{(3+)p}}$ (designated as $C_{2}^{\text{(3+)p}}$) by means of our statistical extrapolation approach. These coefficients normalized to the coefficients $C_{2}^{\text{Mp}}$ for the corresponding many-potential contributions are shown in Fig. 7(b). These data partially explain the convergence-acceleration mechanism. In the area of low $Z$, the coefficients $C_{2}$ for the (quasi-)three-plus-potential contributions decrease by one or two orders of magnitude compared to their counterparts for the many-potential contributions. In the region of high $Z$, they also decrease, but less effectively. It should be noted that for each gauge the reduction of the coefficient $C_{2}$ for both studied acceleration schemes is almost identical, which confirms that the SC trick indeed provide a good approximation to the two-potential contribution.
Figure 8: The individual contributions $\Delta\varepsilon_{a}^{\text{(3+)p}}(k)$ and $\Delta\tilde{\varepsilon}_{a}^{\text{(3+)p}}(k)$ normalized to the total SE correction and multiplied by the factor of $k^{3}$: (a) for $Z=18$ argon; (b) for $Z=92$ uranium.
The subtraction of the two-potential or quasi-two-potential contributions from the many-potential contribution apparently does not change the leading asymptotic behavior of the PW-expansion terms for the remainders $\Delta\varepsilon^{\text{(3+)p}}_{a}$ and $\Delta\tilde{\varepsilon}^{\text{(3+)p}}_{a}$ and only decreases the absolute values of the corresponding coefficients. To illustrate this, in Fig. 8 we plot the individual terms of the PW expansions for these remainders normalized to the total SE correction and additionally multiplied by $k^{3}$, exactly as it was done in Fig. 6 for the many-potential contributions. One can see that the individual terms tend to constants in all the cases, i.e. $p=3$ for (quasi-)three-plus potential terms. That statement has also been verified empirically: the extrapolation procedure gives obviously incorrect results if one omits the term $C_{2}/k^{2}$ in the Eq. (47). We stress that our conclusion contradicts the one from work Sapirstein and Cheng (2023). Fig. 8 also demonstrates that for both acceleration schemes the approach to the asymptotics in the Coulomb gauge is much faster than in the Feynman gauge. The last statement makes it easier to extrapolate the corresponding contributions and certainly makes the Coulomb gauge more preferable for SE calculations within the convergence-acceleration schemes for a wide range of $Z$.
@@ IV.3 Data: summary
In this section, we present and discuss the results of our calculations of SE corrections to the $n=1$ and $n=2$ energy levels in hydrogen-like ions.
First, in Table 2 we study the SE correction for the ground $1S_{1/2}$ state of hydrogen-like argon ($Z=18$), xenon ($Z=54$), and uranium ($Z=92$). As was noted above, Ref. Hedendahl and Holmberg (2012) is considered to be the benchmark in this case. Therefore, in this part of the calculations, we adopt the values of the fundamental constants and nuclear parameters to be the same as those used in that work. For each ion, we consider two gauges, the Coulomb gauge and the Feynman gauge, which are indicated in the second column as ‘‘C’’ and ‘‘F’’, respectively. The corresponding zero- and one-potential contributions, calculated in momentum space, are shown in the third and forth columns. For each gauge we compare different numerical approaches to the evaluation of the the many-potential contribution $\Delta\varepsilon_{a}^{\text{Mp}}$: (i) the standard method, labeled in the column ‘‘Scheme’’ as ‘‘DIR’’, in which the many-potential contribution is treated as a whole; (ii) the two-potential acceleration scheme, labeled as ‘‘TP’’, where this contribution is represented as the sum $\Delta\varepsilon_{a}^{\text{2p}}+\Delta\varepsilon_{a}^{\text{(3+)p}}$, and (iii) the Sapirstein-Cheng acceleration scheme, labeled as ‘‘SC’’, where it is given by the sum $\Delta\tilde{\varepsilon}_{a}^{\text{2p}}+\Delta\tilde{\varepsilon}_{a}^{\text{(3+)p}}$. We also compare two methods for representing the electron propagators: the finite-basis-set method within the DKB basis and the Green’s function approach, which are indicated as ‘‘DKB’’ and ‘‘GF’’, respectively. All the coordinate-space contributions, $\Delta\varepsilon_{a}^{\text{Mp}}$, $\Delta\varepsilon_{a}^{\text{2p}}$, $\Delta\varepsilon_{a}^{\text{(3+)p}}$, and $\Delta\tilde{\varepsilon}_{a}^{\text{(3+)p}}$, are obtained by applying the statistical extrapolation procedure described in Sec. III.3. Within the DKB approach, the PW-expansion series are truncated at $|\kappa_{\text{max}}|=20$. When using the GF approach, these series are truncated at $|\kappa_{\text{max}}|\sim 45$ in the case of the two-potential contribution $\Delta\varepsilon_{a}^{\text{2p}}$ and at $|\kappa_{\text{max}}|=24$ in the other cases. Finally, the total SE corrections obtained by means of different approaches are given in the last column. For each ion and gauge, the results from Ref. Hedendahl and Holmberg (2012) are shown for comparison as well.
As can be seen from Table 2, the main source of numerical errors, regardless of the gauge, is the many-potential contribution. Even when applying the acceleration schemes, this contributions still limits the accuracy of the total SE corrections. We note that the many-potential contribution is significantly smaller in the Coulomb gauge than in the Feynman gauge (see the penultimate column in Table 2), which was also noticed in Refs. Hedendahl and Holmberg (2012); Yerokhin et al. (2025a). However, for the reasons described above, this does not provide a significant increase in accuracy unless acceleration schemes are involved. Although obtaining the zero- and one-potential contributions with sufficient accuracy is not a technically difficult task, an attractive feature of the Coulomb gauge, as can be seen from Table 2, is the absence of large cancellations between these momentum-space contributions. Nevertheless, this observation is extremely useful for obtaining high-precision results for low-$Z$ systems.
Clearly, acceleration tricks reduce extrapolation uncertainties and allow for a more accurate evaluation of the many-potential contributions. Using either the two-potential scheme or the SC scheme reduces the numerical errors and allows one to obtain up to two additional significant digits. As a rule, the most accurate many-potential contribution can be obtained by using a combination of the Coulomb gauge and the SC acceleration scheme.
The GF approach provides a more accurate treatment of the many-potential contribution compared to the DKB approach. This results mainly not from the lower value of $|\kappa_{\text{max}}|$ employed in this case, but from the fact that the DKB method needs the approximation with respect to the size of the basis set. The low accuracy of the DKB approach does not detract from the already mentioned advantages, namely, the ability to exclude a specific state and construct Green’s functions for a wider class of systems that do not possess spherical symmetry. In some sense, the FBS method (not necessarily in the DKB realization) is more flexible and easier to implement. Therefore, if high accuracy is not required, it is an excellent method for calculating the SE correction.
Returning to the comparison with the benchmark calculations Hedendahl and Holmberg (2012), we note that our results obtained by means of different methods are in good agreement with each other and those from Ref. Hedendahl and Holmberg (2012). Small deviations have been revealed. However, the reason for them is unclear for us.
For a broader comparison, another set of calculations has been performed. Namely, we consider the SE correction for the ground $1S_{1/2}$ state in hydrogen-like ions with $Z\in\{10,18,26,36,54,82,92,100\}$. In this case, we compare our results with those from Ref. Yerokhin and Shabaev (2015a). Therefore, the nuclear parameters are taken from Ref. Yerokhin and Shabaev (2015a). Average values of rms nuclear radii from there were used. An estimation of the uncertainty of the SE correction with respect to the nuclear radius and nuclear model is beyond the scope of the current paper, the interested reader can find it in the original work Yerokhin and Shabaev (2015a). The values of fundamental constants, however, have been adopted from Ref. Tiesinga et al. (2021). The corresponding results are summarized in the Table 3, which is organized similarly to the Table 2. Overall, the results shown in Table 3 confirm the findings obtained earlier. The relative smallness of the Coulomb-gauge many-potential contribution is illustrated by Fig. 9(a), where different contributions to the SE correction for the $1S_{1/2}$ state in both gauges are plotted as functions of $Z$. This smallness as well as the absence of cancellation between the momentum-space contributions in the Coulomb gauge is demonstrated also in Fig. 9(b), where the sum of the zero- and one-potential contributions is shown. For neon ($Z=10$), a more ‘‘regular’’ behavior of the zero- and one-potential contributions in the Coulomb gauge is even more pronounced than for higher-$Z$ systems. From Table 3, it is seen that our results are in reasonable agreement with the results from Ref. Yerokhin and Shabaev (2015a). We conclude that the most accurate calculations of the many-potential contributions can indeed be performed using the combination of the Coulomb gauge and the SC convergence-acceleration scheme.
Using the most effective methods among those examined, namely, the GF approach and the SC convergence-acceleration scheme, we consider the SE corrections for the $n=2$ states: $2S_{1/2}$, $2P_{1/2}$, and $2P_{3/2}$. The SE corrections are calculated for hydrogen-like ions with $Z\in\{10,18,26,36,54,82,92,100\}$. The corresponding results for both Feynman and Coulomb gauges are given in Table LABEL:tab:we_yerokh_excited_states. Analyzing the table one can see that the accuracy of the total SE corrections for the excited states in hydrogen-like ions is still limited by the uncertainty of the many-potential contribution calculations. As in the case of the ground state, this contribution in the Coulomb gauge is significantly smaller than in the Feynman gauge. Moreover, the Coulomb gauge as before lacks strong cancellations between the momentum-space contributions. Finally, we note that our results are found to be in good agreement with the ones from Ref. Yerokhin and Shabaev (2015a).
@@ V Conclusion
The self-energy correction for various states of hydrogen-like ions is considered. Two gauges, Feynman and Coulomb ones, are studied, the acceleration tricks are discussed, and a self-consistent set of useful for the corresponding calculations formulas are collected.
The acceleration schemes do improve the accuracy of calculations. The combination of the Coulomb gauge and the Sapirstein-Cheng partial-wave-expansion convergence-acceleration scheme Sapirstein and Cheng (2023) is most suitable (among those considered) for calculating the self-energy correction in hydrogen-like ions, especially when studying light nuclei.
We believe that the analysis performed in the present work will facilitate the calculation of a wide range of more complex radiative corrections with the self-energy loop.